OpenStudy (anonymous):

You win $4 if you flip a coin twice and get tails both times. For any other outcome you get nothing. What is the expected value?

OpenStudy (anonymous):

expected value in this case is the amount you win times the probability you win it

OpenStudy (anonymous):

the probability you get tails both times is \[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\] and \[\$4\times \frac{1}{4}=\$1\]